Joseph B. Eastman— A Study in Public Service

Claude Moore Fuess^{∗}

This paper is a somewhat hazardous venture into almost contemporary biography—hazardous for at least two reasons. Many of those in this room, especially those with railroad connections, may have known “Joe” Eastman very well and may have had dealings with him of which I have heard nothing. On the other hand, because he was never a publicity seeker, like General Hugh Johnson or Harold Ickes, some of you may confuse him with his distant and more opulent cousin, George Eastman, the Kodak King. Thus, for my present purposes, some of you may know him too well and others not well enough. I may possibly learn more from you than you will from me; but my mission will be discharged if I can at least make clear what he was and what he did.

During the Hoover administration, when Joseph Bartlett Eastman had been for more than a decade on the Interstate Commerce Commission, three of us were lunching with him, and somebody said to Joe—everybody called him “Joe”—“I hear you’ve been offered a railroad job at a salary of more than $50,000.” “It could be,” he cautiously admitted. “It isn’t your first offer either, is it, Joe?” “Maybe not,” was the modest answer, in the Amherst tradition. Then I put in my oar—he had been a year before me at college— and asked, “Aren’t you ever tempted to accept a position like that, with an enormous income, as compared to what you get slaving for the government, and with a lot less grief?” “Well,” replied Joe meditatively, “probably it is a luxury for me to stay on here, just as it would be to own a yacht, but I guess I can afford it.”

This reply was characteristic of the person of whom Mr. Justice Brandeis once said, “Joe Eastman has more interest in the public service and less in his own career than any man I have ever known.” From the day he was graduated until he died almost forty years later, Joseph B. Eastman was constantly working for the people, not for management or for labor, not for railroad operators or railroad employees or shippers, but for the good of the American economic and social order. Although he never ran for office and was even in doubt concerning his political party, he held one appointive job after another in gradually widening areas. Calvin Coolidge never held an appointive office; Eastman never held an elective one. When he died at the age of sixty-two, worn out like an over-used internal combustion engine, it was said of him that he had “set a pattern of intelligent devotion to the public welfare which, if extensively recognized, would greatly uplift the character of public service in the United States.”

∗This paper was read at the January, 1950, meeting.

One incident will show the quality of his integrity. On December 19, 1918, President Wilson nominated Eastman to the Interstate Commerce Commission—at only thirty-six, he was the youngest man ever named for that office. The Commission legally had to be bipartisan, and the membership was such that the existing vacancy could be filled only by a Republican. When Eastman heard the news, eager though he was for the position, he promptly wrote to Senator Lodge to explain that he was an independent in politics and could not permit a misleading label to be attached to him. No one in the Senate chose to make an issue of the matter, and he was confirmed without any discussion; he never knew exactly why or how. Again and again, Eastman had to explain, to the embarrassment of his sponsors, that he had no party affiliations. As a matter of fact, he voted for Taft in 1908, for Theodore Roosevelt in 1912, for Wilson in 1916, for Cox in 1920, for La Follette in 1924, for “Al” Smith in 1928, for Franklin D. Roosevelt in 1932 and 1936, and for Willkie in 1940. No one could have been more discriminating, or more completely independent.

His independence was supported by equally conspicuous courage. When he was confirmed for the Interstate Commerce Commission on January 24, 1919, the railroads were still under wartime operation by the federal government; and Director-General McAdoo had formally recommended that national control should be extended for a period of five years, until January 1, 1924. Soon after, the Commission, with one member dissenting, presented to the Senate its official conclusion that “with the adoption of appropriate provisions and safeguards for regulation under private ownership, it would not be wise or best at this time to assume Government ownership or operation of the railroads of this country.”

When Eastman took the oath of office, then, he was aware that the Commission had placed itself on record in favor of a return to private operation and that it was unnecessary at that moment for him to express his personal views. Nevertheless, after much sleepless soul-searching, he sent direct to the Senate Committee on Interstate Commerce a comprehensive letter arguing for the continued operation of railroads by the federal government and maintaining that, on the whole, it had been a success during the war. He reached this conclusion, not impulsively, but by a process of measured reasoning on the basis of the evidence, as Robinson Crusoe balanced his blessings against his calamities. George Creel once remarked to Eastman’s sister that Joe’s thinking reminded him of an elephant’s crossing a bridge—putting down each foot carefully and testing the strength of the structure before making an advance. When she told her brother of Creel’s comment, he smiled and replied, “Ask him if he ever heard of an elephant’s falling through a bridge. He usually somehow gets across the stream!”

No small amount of temerity was required for the youngest member of the Commission in both age and service to take issue with his new colleagues on such a fundamental and controversial matter. Some of them were annoyed, and in private called him “fresh” and “naive,” but they shortly had to recognize that they had in their midst that rare bird, a completely uncontrolled thinker—a man who cared more for the approval of his conscience than for his personal advancement. Washington lobbyists discovered that Eastman could not be bribed or flattered or coerced. Somebody said to me last spring, “Joe Eastman set the Interstate Commerce Commission in a turmoil shortly after he arrived in Washington and kept it boiling until he died.” Although it wasn’t quite so bad as that, he did become on the Commission what Mr. Justice Holmes had already become on the Supreme Bench—the member whose dissenting opinions were closely studied and universally respected.

Different types of Presidents would have been glad not to reappoint him— Harding in 1922 and Hoover in 1929, because he was too liberal, and Roosevelt in 1936, because he was too conservative. But when the choice had to be made, they did not dare to face the outcry which would have arisen if they had turned him down. During his long career, Eastman became hardened to abuse. Until 1933, railroad executives felt that he was a radical, an amateur who did not understand their difficulties. After 1933, he was condemned by labor leaders as a reactionary because he would not support their demands. I cannot see that he was prejudiced at that period. He was serving as best he could the interests of the general public, let the chips fall where they might.

How did he develop his political and economic philosophy? He was born in 1882, in the village of Katonah, New York (now inundated, although for the moment rather feebly, by Manhattan’s water supply), where his father was the Presbyterian clergyman. He came from sound Yankee stock and an exceptionally literate home. In Amherst, which he entered from Pottsville, Pennsylvania, his father’s second parish, he was distinguished chiefly as editor of the college paper and leader of the debating team. His philosophy professor, the eminent Charles E. Garman, was accustomed to lead his students through the morasses of atheism and agnosticism to the safe meadow of orthodoxy, following a trail which he had carefully charted. Joe Eastman dared to press him with embarrassing queries which interrupted the teacher’s progress. Finally the ailing and impatient professor sent for Joe and said, “Mr. Eastman, your skepticism is ruining my class. Will you kindly refrain from further disputation or else cease to attend it!” After this rebuke, Joe sat silently through the remaining lectures, but his comments in the fraternity house were pungent and resentful. Later, from the bench of the Interstate Commerce Commission, Eastman sometimes punctured attorneys’ arguments with the same type of penetrating, irritating questions; but they could not silence him, as Professor Garman had done.

Eastman’s first intercollegiate debate, with Bowdoin, found him arguing the negative of the subject, “It is for the public interest that employers recognize trade unions in the arrangement of wage schedules”—and, believe it or not, he carried his team to victory! Then, as later, an incurable misogynist, he presented eloquently the affirmative of the question, “Resolved, that the proximity of Smith and Mount Holyoke is detrimental to the interests of Amherst College, in the broad sense.”

In public speaking, as in everything else, Eastman was a slow starter but a good finisher. For example, although he was far from being a natural athlete, he made himself into an excellent and most exasperating tennis player, who often, by his “lobs,” tempted his opponents into recklessness, and defeated competitors rated as better than he. He was elected to Phi Beta Kappa, but only at the second drawing. He was not very popular as a Freshman, but ended as Class President in his Senior year. Nobody ever described him as “brilliant.” The adjectives applied to him by his classmates were “steady” and “reliable,” “dependable old hoss”—and it was the same all the rest of his life.

The chief influence on his thinking as an undergraduate was not the faculty, but President Theodore Roosevelt, the “fighting champion of the people.” Eastman and I were graduated in that antediluvian Golden Era between the Spanish War and World War I—in America, at least, probably the happiest and most hopeful epoch this planet has ever known. However it may have been at Harvard in those days, up in the Connecticut Valley we had no doubt that the millennium couldn’t be very far off. The idea of progress through evolution and the doctrine of human perfectibility were in the air, and we had no Spengler or Toynbee to disillusion us. Nobody could possibly have predicted, or have found any reason for predicting, the tragic years that lay ahead. “Teddy” Roosevelt, the crusader, the leader of reforming movements, was for Joe Eastman a knight in shining silver, the warrior who every man-at-arms would wish to be.

Eastman apparently never thought of going into business, like many of his classmates, nor did his religious views permit of his becoming a clergyman, like his father and grandfather. His desire to help the underprivileged in a practical way led him to accept a South End House Fellowship of five hundred dollars, in Boston, under another fascinating personality—the late Robert A. Woods. Union Park, only a few blocks from the Ritz-Carlton and Bonwit Teller, was, and is, a district of boardinghouses and cheap apartments, where Woods, also an Amherst graduate, was carrying out a program of “social reconstruction.” There Eastman settled in the autumn of 1904, to learn how the other half lived. Woods was a truly remarkable philanthropist, a pioneer in “settlement work,” with a gift for attracting younger men into his orbit and imbuing them with his indomitable, altruistic spirit. Under his direction and inspiration, Eastman was soon conducting classes, listening to hard-luck stories, and trying in his inexperienced way to alleviate human misery.

But it was the operation of government, and particularly of municipal politics, that attracted him most. Woods tried intelligently to “size up” his assistants and guide them into fields where they would be most effective— and when he ascertained Eastman’s bent, he encouraged him to accept the secretaryship of the Public Franchise League, a small informal group of highminded citizens backed by a Boston lawyer, Louis D. Brandeis, and a Boston merchant, Edward A. Filene. Brandeis, seeing the potentialities of the young Eastman, made him his confidential assistant in his battles with the Boston Elevated, the Consolidated Gas Company, and the New York, New Haven and Hartford Railroad. Years later, Eastman wrote, “I was a boy of New England stock and conservative tendencies. The association with the men in the League, and particularly Justice Brandeis, gave me a new point of view and undoubtedly changed the current of my life. Justice Brandeis had the most powerful intellect with which I have ever had the good fortune to come in contact, and I have met a goodly number of eminent men.”

Under Brandeis’s sponsorship, Eastman soon had a tiny office in the Equitable Building, with a salary of $1,000 (usually paid in small weekly checks from Mr. Filene), an ancient typewriter, and plenty of pigeons to feed on the window sill. He soon became acquainted with the play of sinister personal and social forces beneath the surface of boss rule and ward politics. He learned much about the practical administration of the city government. He also became involved in Brandeis’s long duel with Charles S. Mellen, President of the New Haven and the Boston and Maine railroads, and trained himself to read a balance sheet with an appraising eye. He carried on research projects, made digests of documents, appeared before legislative committees; indeed, he assembled most of the statistics for the famous Interstate Commerce Commission Report which described the record of the New Haven as “an amazing narrative of reckless and profligate financial management.”

Theodore Roosevelt, Robert A. Woods, Louis D. Brandeis, all three, concurrently and successively, each in a different way, moulded the thinking of the sensitive and susceptible Amherst graduate; but it was Brandeis whose character, ideas, and methods left upon him the deepest impression. By 1913, however, Mellen had retired under fire, and Eastman, urged by his friends and feeling that his job for the Public Franchise League was completed, resigned as its secretary, planning to resume some legal studies which he had undertaken years before. Meanwhile, the employees of the Boston Elevated Railroad had formed a carmen’s union and instituted proceedings for higher wages. The union leader, James H. Vahey, asked Brandeis to recommend somebody competent to act as its counsel. “I will name a man for you if you will let me fix his compensation,” answered Brandeis. “Why do you want to do that?” inquired Vahey. “Because he doesn’t know his own value well enough to fix a proper compensation for himself!”

So it was that Joseph B. Eastman, although not an attorney, accepted the union’s retainer and entered at once on a comprehensive investigation of the wage system of the Boston Elevated. He had learned how to uncover, assimilate, and integrate facts; and he put up such persuasive pleas before the arbitrators—James J. Storrow, James L. Richards, and James H. Vahey— that they were unanimous in recommending a comprehensive increase in pay for the men. Both Storrow and Richards, although more conservative than Eastman, were tremendously impressed by his arguments and became his friends for life. A few months later, as counsel for another union which was pressing the Middlesex and Boston Street Railway Company for higher wages, Eastman moved from the immediate issue into the broader field of economics, maintaining that the financial condition of the company should be immaterial to the question of fixing fair compensation, and that wages, in whatever circumstances, should be sufficient to enable a workman to support his family in health and comfort, and to build up adequate reserves for his old age. His reasoning undoubtedly had a powerful effect on the policies of public-service corporations.

Eastman at this period was guaranteed election if he would run for the General Court, but his experience on Beacon Hill with what he called “the small army of professional lobbyists” had discouraged him. To a friend he wrote, “It is a crime the way this Legislature is pulled and hauled by private interests.” For the moment he was uncertain of his future. He had gained immensely in self-confidence, but no attractive opportunity presented itself. In 1913, however, the Washburn Bill had met a popular demand by creating the Public Service Commission of Massachusetts; and a year later, when one of its original members, George W. Anderson, resigned, Governor David I. Walsh disregarded the warnings of his political advisers and appointed Eastman to the Commission, at a salary of eight thousand dollars. He was then only thirty-three. One of his fraternity mates wrote him, “While I don’t agree with a lot of your crazy ideas I will do you the justice of saying that I believe you are sincerely honest and will be open to conviction, and that is why I am glad to see you appointed.” This was a feeling which was shared by many others throughout the Commonwealth.

Before Eastman had been on the Public Service Commission six months, he was recognized as its dominating and best-informed member. Soon he had to confront the Bay State Rate Case, the most complicated of its kind that any Massachusetts commission had ever been called upon to settle; and he, solus et unus, prepared the report and order requiring the company to cancel a proposed new schedule of increased charges. It was in this report that he committed himself to the “prudent-investment” theory of valuation of which he was later to become the ardent exponent. He also drafted a careful analysis, covering 317 pages, of the financial situation of the New Haven Railroad, reaching the conclusion that the monetary transactions of that company had been “numerous, intricate, and confusing”—certainly a masterpiece of understatement. Never until World War II did Eastman reach the stage where he trusted the management of the New Haven; and his frequent and uninhibited correspondence with Mr. E. G. Buckland, its vicepresident and general counsel, is packed with criticism, satire, and some very blunt moralizing.

By this time, however, the Boston Chamber of Commerce had decided that Eastman did not really have horns and a tail, and that he might be justified in some of his accusations. His good angel, Brandeis, although vigorously opposed by a group of fifty-five of the “solid men of Boston,” had been approved in 1916 as President Wilson’s nominee for the Supreme Court, and had moved to Washington. Two years later, when a vacancy unexpectedly occurred on the Interstate Commerce Commission, Brandeis recommended Eastman and, to secure corroboration for his own admittedly partial opinion, asked Commissioner Robert W. Woolley to go to Boston and look Eastman over. Woolley, a Kentucky Democrat, had no idea what Brandeis had in mind, and Eastman himself had only vague suspicions. Although Woolley was eleven years older, the two men took to each other spontaneously, and Woolley reported to the President, “That man Eastman is one of the most remarkable fellows I ever met.” This, of course, was no news to Brandeis.

Eastman’s professional qualifications could have been catalogued as follows: a considerable familiarity with transportation facts and figures, and a broad knowledge of common carriers and their problems; an intimate acquaintance with the difficulties of employees, gained while he had been counsel for the unions; a theory, already well formed, regarding the proper relationship of transportation, in all its phases, to the communities which it serves; an orderly economic philosophy, the product of practical experience and much meditation; and a devotion, deep-seated and unselfish, to the public welfare. However young he was in years, no other member of the Commission had ever been better equipped to meet its peculiar responsibilities.

Eastman was now, at thirty-seven, identified with what has been called by one authority, with some obvious exaggeration, “the most powerful body in the world, administrative or otherwise.” At the close of the report on the very first case in which he participated are the words, “Eastman, Commissioner, dissents”; and from that date on we find him announcing, “I find it difficult to go along with my colleagues,” or, “I am in doubt in regard to this case,” or, “I must frankly express disagreement.” Any honest man’s response to current issues is bound to reflect his basic creed. Eastman’s opinions, both concurring and dissenting, were the inevitable product of firm convictions, and as a group constitute a pattern the elements of which blend into a coherent and orderly whole. Even his superficial inconsistencies prove on examination to be reconcilable. Any one familiar with his habits of thought and his estimate of relative values could usually predict his reaction to completely new problems. Unforeseen considerations naturally had to be adjusted to his principles. Occasionally the arguments were so much confused as to create doubt. But Eastman was always, consciously or unconsciously, aiming at justice for Mr. Average Citizen; and his reflections over the years led him to conclusions which fitted together like iron filings clustering around a magnet. He never dissented from himself. Once he said in Washington, “They’ve been calling me a ‘crack-pot,’ but I guess, after what I’ve gone through, I can take it. Even if I am cracked, my theories will hold water!” Dissenting opinions on the Interstate Commerce Commission accomplish almost nothing, except to enable “Lone Wolves” to howl. The decision of the majority, whether in a division or in the full Commission, even when close, is final; and Eastman’s vigorous expressions of disapproval had no effect on the fortunes of the litigants. Over a quarter of a century, however, he did accomplish much in the education of his colleagues as well as of himself.

It is not my intention to inflict even on the tolerant members of this Society any dissection of the technical questions with which Eastman and his colleagues had to struggle. To a layman the reports of the Commission often seem to be written in a “maze of jargon.” What is a “combination rate” which “makes on” the Ohio River? What is “fourth section relief”? What is a “joint through rate”? What is “Rule 77”? What is the “intermediate rule”? As I have perused diligently the records of the Commission, I have found myself learning what amounts to a new dialect, if not an actual foreign tongue. Considering his lack of legal training, Eastman mastered this lingo with amazing rapidity, and very few suspected that he had never taken bar examinations.

However puzzling their language, his opinions had one common denominator—a regard for the public good. The “recapture clause” of the Transportation Act of 1920 made it necessary for Eastman to pay much attention to the valuation of railroads—an unromantic and never-ending task. His advocacy of the “prudent-investment” theory as contrasted with that of the “cost of reproduction” was opposed by the railroads, by several of his Commission associates, and even by the Supreme Court; but he stuck by his guns, believing that he was working for the good of the people, and his position has been vindicated. He made mighty efforts to reduce the costs of operation by consolidating competing lines; but he was opposed to the formation of holding companies to bring about combinations for which authorization could be obtained through no other process. He did not like banker control of railroads and sharply condemned the conduct of certain companies in allowing their securities to be marketed by single preferred banking houses. He criticized the depletion of railroad resources by what he described as “opulent awards made to bankers and lawyers in connection with reorganization.” In the 1930’s, when railroad revenues were falling off and wages had been reduced, he urged railroad executives to relinquish part of their salaries voluntarily and actually succeeded in getting them lowered to a maximum of $60,000 a year. In sugaring this bitter pill, Eastman remarked:

One thing certain is that money is by no means the only compensation received by a railroad president or even by a lesser executive. The best compensation of all—in my judgment a more effective one than is commonly supposed—is the joy of creative work well done, particularly when it involves the element of public service.

Mr. Thomas W. Lamont, of Morgan and Company, once expressed this high ideal in almost precisely the same words—a fact all the more interesting because Lamont, like Eastman, was a minister’s son, brought up, like Eastman, in a Katonah parsonage.

The Great Depression which overwhelmed the country after 1929 had, of course, a disastrous effect on the railroads, which had not done badly in the 1920’s. Some relief was essential; and early in the first Roosevelt administration, following suggestions offered by Eastman, the Emergency Railroad Transportation Act was passed by the Congress, and Eastman was shortly appointed by the President as Federal Coördinator of Transportation. One primary purpose of the measure was to eliminate waste; but unfortunately the railroad labor organizations tied Eastman hand and foot by insisting on the provision that reforms must not further reduce the number of employees. Eastman, once again with the interests of the public in mind, opposed this clause, saying, “I do not believe the way to solve the problem of unemployment is to retain work which is lost motion, and amounts to waste and inefficiency.” In his attempt to control “feather-bedding,” Eastman was blocked by the unions, which did all that they could secretly to undermine his authority. He now found himself in the unhappy position of being upbraided by his former allies; and some union leaders, completely lacking in gratitude, never forgave him for not going all out for their cause.

Although in his first report as federal coördinator Eastman reiterated his conviction that, as he phrased it, “theoretically and logically public ownership meets the known ills of the present situation better than any other remedy,” he was not ready to advise it in 1934. He had always regarded the question as one of practical expediency rather than of political theory, and he now felt that theory and logic could be disregarded. It was no time, he stated, for the federal government to assume several billion dollars of additional debt; and he was sure that under federal control very few savings could be accomplished if the unions persisted in their obstinacy. When he announced his views, conservatives hailed him as a sinner returned to the fold.

Eastman tried desperately as federal coördinator to live up to his title and please everybody, but succeeded in placating neither management nor labor. The companies promptly reorganized the almost defunct Association of American Railroads and through it did all they could to thwart some of his most cherished plans. In commenting on their attitude, Eastman said:

At bottom, the trouble is that the managements think narrowly in terms of their own particular roads, rather than in broader terms. Quite naturally and quite properly they put the welfare of their own particular road first and foremost. What they do not appreciate, as I think they should, is the extent to which the individual welfare will be promoted by action which is for joint benefit or common good.

Once Eastman burst out, somewhat ingenuously, “I never realized until I took this job how many completely selfish pressure groups could mess things up in Washington.” Even when he produced a highly intelligent plan for reorganizing the Interstate Commerce Commission, a few of his colleagues, apparently jealous of his power, managed to frustrate his designs. By sheer persistence, however, he was able to push through Congress measures for the regulation of common carriers on the highways and of water carriers in interstate and foreign commerce. It was at this time that he declared, “The transportation system is a unit and must be dealt with as such. … The various agencies interlock and react, one against another, in a multitude of ways. The system cannot permanently be half regulated and half unregulated. If the principles of a battle royal are to govern, it is unfair to handcuff the railroads.”

The office of Federal Coördinator of Transportation lapsed in 1936, without many regrets, and Eastman returned to his regular duties as a member of the Interstate Commerce Commission. During his three years he had initiated and carried through an unprecedented amount of special research, and more than sixty separate reports, packed with information, were issued from his office. In addition, he delivered fifty-two prepared addresses and numerous informal talks. It is true that although in theory he possessed great power, he thought it injudicious to exercise it, but he did set the tone of thinking for the railroad world. One eminent authority, commenting on Eastman’s achievements, has remarked, “His ideas, although but slowly accepted, will inevitably lead to improvements on a plane never before directly reached by our government.”

Although Eastman respected President Roosevelt, he found him difficult to deal with, especially on matters in which organized labor had a part; and he was hardly on the periphery of the White House inner circle. On one painful occasion the President, having summoned to his office the heads of various important government agencies, asked them to send him direct their reports in advance of publication. One by one those present agreed, some of them reluctantly; but when Eastman’s turn came, he said, “I’m sorry, Mr. President, but as you must be aware, the Interstate Commerce Commission is an agent of Congress and created by it, and it is my duty to report to it first.” Mr. Roosevelt was much displeased by Eastman’s open criticisms of labor officials and in reprisal, when the latter’s term expired on December 31, 1936, he did not reappoint him until July 9, 1937. In the interval, of course, Eastman continued to function, although uncertain of his status. He strenuously objected to having the independent regulatory commissions put in a position where the performance of their function would be subject to executive control. He trusted then, and always, in the efficiency of small governing bodies properly organized and devoted to the public interest. As federal coördinator, he had had his own chance to become a dictator, and had rejected it.

Eastman was now conceded to be the foremost authority in the country— indeed in the world—on transportation, and was consulted by the Congress on all legislation relating to common carriers. In 1939, his fellow members of the Interstate Commerce Commission showed their confidence in him by voluntarily choosing him as chairman for three years instead of for the customary one-year term. When we entered World War II in December, 1941, it was inevitable that everybody, including the President, should turn to him for leadership in the field of which he was the acknowledged master. As Director of Defense Transportation, charged with the coordination of shipping on railroads, motor carriers, and water carriers, he took a prompt stand against government operation, thereby securing the fullest cooperation of railroad executives. They unquestionably did their best in the emergency, and were even disposed to alter their regulations upon request without the necessity of coercion. Eastman influenced transportation agencies as much as possible through education and persuasion and was greatly pleased with the results. Indeed, he publicly attributed the excellent performance of railroad operation during the war to the decision to leave the management in private hands. Again his criterion for measuring efficiency was the result on the public welfare.

On February 17, 1944, when a group of his admirers held a dinner in Washington to celebrate his completion of a quarter of a century of service on the Interstate Commerce Commission, Eastman responded in what he called a “twelve-point primer,” summing up what he had learned over the years. Speaking of the Commission and similar bureaus, he asserted emphatically that “they must not be under the domination of either the President or Congress or of anything else than their independent judgment of the facts and the law.” In one illuminating confessional paragraph, he said:

Sitting in dignity and looking down on the applicants from the elevation of a judicial bench has its dangers. A reversal of the position now and then is good for the soul. It has for many years been my good fortune to appear rather frequently before legislative or congressional committees. They are a better safeguard against inflation than the O.P.A.

Commenting reminiscently on the attacks which he had been obliged to endure from vested interests, he said:

Power is not a permanent but a shifting thing. I can well remember the time when it was dangerous to incur the displeasure of bankers, but there has been no danger in this since 1932. It later became a great danger to incur the displeasure of farm or labor organizations. There is nothing more important than to curb abuse of power, wherever it may reside, and power is always subject to abuse.

A few weeks later “Joe” Eastman was dead at sixty-two, the victim of unceasing toil and cumulative fatigue. Then, as so often, he reaped his reward. His passing at that critical moment in the war was viewed as a national calamity. Even those who had disagreed with him testified to his integrity and exceptionally high ideals, and he was hailed by conservatives and liberals alike as the perfect public servant.

Some of you must have seen him sitting with his elbows resting on the table in that imposing room reserved for Commission hearings. By middle life he had become rotund, with rosy cheeks and a domed forehead from which the brown hair was slowly receding. His manner was affable and he seemed to be easygoing, but his jaw was square and firm, and his blue eyes could be keen and searching. Until he was almost sixty, he kept himself in condition by playing squash and tennis, and he usually spent his summer vacations in Canada indulging in that sport of gentlemen—dry-fly fishing. Aside from these he had few recreations. A decidedly eligible bachelor, he seldom went out to social affairs, and a hostess who secured him for a cocktail party was justified in boasting of her prize. It would be wrong to imagine him as discouragingly virtuous. Although he was temperate in his habits, he was no ascetic and knew well the difference between Scotch and Rye. His collection of pipes was the envy of all his friends.

For years Eastman lived with his sister in a modest house in Cathedral Road, walking each morning to his office, a distance of nearly three miles. He never owned an automobile, even when his salary would have justified it, preferring to travel by taxicab. He was a simple, unassuming person who detested “stuffed shirts.” In 1943, he escorted a lady to the Union Station and climbed aboard her train to see her off. Somehow it started before he could dismount. When the conductor came along and found that Eastman had no ticket, he said, “That’s just like a lot of you fellows—don’t buy a ticket and then think you can get away with it!” Joe made no reply, but quietly paid his fare. As the conductor continued down the aisle, still sputtering, somebody said to him, “Don’t you realize who that is?” “No— and I don’t give a damn!” was the reply. “Well, that’s Eastman, the Director of Transportation.” “Joe Eastman? My God!” cried the conductor, and then started back to apologize. Joe wouldn’t even let him open his mouth. “You’re absolutely right, conductor,” he said, “and I was entirely to blame.”

Eastman’s passion and capacity for continuous hard work made him a legendary figure in the Interstate Commerce Building. Evening after evening he would stay alone in his office, sitting in a deep chair with a board across his knee, pondering columns of figures. When he had charge in the 1920’s of the notorious Southern Class Rate Investigation, his two ablest assistants unexpectedly resigned, and he carried on virtually alone the hearings in what was probably the most complicated matter ever to come before the Commission. He called his report on this case his magnum opus; and indeed it was to him what Paradise Lost was to Milton, or The Forsyte Saga was to Galsworthy. Much of this labor was unnecessary, but he believed that only through a personal study of the issues involved could he develop a proper understanding of the problems and be just to everybody.

With this desire to be fair-minded went a genuine kindness of heart and a sympathy for the underdog which went back to his early days in South End House. One of his weaknesses was a disposition to judge subordinates too leniently, and he had the reputation of being gullible and too softhearted. It was not difficult to hoodwink once a man whose instinct was to trust those around him. But when he was betrayed or deceived, as occasionally happened, he was terrible in his wrath.

So far as I have been able to determine, Eastman was never tempted by money or power or fame. He once wrote:

It is a common belief that the desire for financial gain is the only motive that will impel men to their best endeavors. I challenge that tenet whole-heartedly. I was brought up in a minister’s family; I have enjoyed the friendship of doctors, schoolteachers, and professors, and I have had an opportunity of observing men in public life as well as many engaged in private business. It is my profound conviction that the best things which have been done in the world have been impelled by higher motives than the desire for gain.

This theory Eastman exemplified in his private life and public career, again and again declining offers which would have made him rich in a short period. When his few needs were satisfied, he asked nothing more in the way of recompense for his services. At least three times he imperiled his reappointment by a free expression of his opinions when the decision was in doubt.

Naturally he had to defend himself against attack. At various periods and by different people he was described as a “radical” and a “reactionary,” “an economic royalist” and a “New Dealer,” but he was never greatly troubled except when he was denounced as “socialistic.” “ ‘Socialistic,’ ” he declared, “is a catchword loosely used as a means of discredit in default of argument.” Writing in 1924 to E. G. Buckland, whose views he did not like, he said:

I thank you for sending me the address of Martin W. Littleton to the American Defense Society. … I have glanced through this address and am not inclined to take much stock in it. It resembles so many of the unwarrantable charges that have become current in the past few years to the effect that various citizens are “reds,” in the pay of Soviet Russia. I know of instances where such charges have been directed against persons with whom I am intimately acquainted and for whom I am able to vouch.

All this has a very modern ring, and I have no doubt that if Eastman were still alive and still on the Commission, some Hickenlooper or McKellar would describe him as un-American. In the end, however, Eastman earned the regard of some of his most outspoken critics. In the 1920’s, when he was assailing the management of the New Haven, one of its officers said to me, “This fellow Eastman is just a bloodhound on our trail—but at any rate he keeps us moving!” Men high up in the councils of the Association of American Railroads have within a few weeks expressed their approval of what he did for their industry. One of them remarked, “He told us the truth, and it was good for us. We knew he had no axe to grind!”